A particle of mass L=M²X^4/12 +MX²V<x>-V²<x>
Where V is some differentiable function of X. Find the equation of motion for X(t) and describe the physical nature of the system on the basis of this equation
1
Expert's answer
2019-09-09T11:14:42-0400

The Lagrangian for the particle with mass "m" is:


"L=\frac{m^2\dot{x}^4}{12}+m\dot{x}^2 V(x)-V^2(x)."

Differentiate by "x":


"\frac{dL}{dx}=m\dot{x}^2\frac{\partial V}{\partial x}-2V(x)\frac{\partial V}{\partial x}.\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space(1)"

Differentiate the last equation by "\dot{x}":

"\frac{\partial L}{\partial\dot{x}}=\frac{1}{3}m^2\dot{x}^3+2m\dot{x}V(x)."

Now differentiate by time:

"\frac{d}{dt}\Big(\frac{\partial L}{\partial\dot{x}}\Big)=m^2\dot{x}^2\ddot{x}+2m\ddot{x}V(x)+2m\dot{x}^2\frac{\partial V}{\partial x}\space\space\space\space\space\space(2)."

The Lagrange's equation:

"\frac{d}{dt}\Big(\frac{\partial L}{\partial\dot{x}}\Big)-\frac{\partial L}{\partial{x}}=0."

Substitute (1) and (2) in the last equation:


"m^2\dot{x}^2\ddot{x}+2m\ddot{x}V(x)+2m\dot{x}^2\frac{\partial V}{\partial x}-\\
-m\dot{x}^2\frac{\partial V}{\partial x}+2V(x)\frac{\partial V}{\partial x}=0,"

"\space\\
m^2\dot{x}^2\ddot{x}+2m\ddot{x}V(x)+m\dot{x}^2\frac{\partial V}{\partial x}+2V(x)\frac{\partial V}{\partial x}=0,\\
\space\\
(m\dot{x}^2+2V)\Big(m\ddot{x}+\frac{\partial V}{\partial x}\Big)=0.\space\space\space\space\space\space\space\space\space\space(3)"

That is the equation of motion.

A force can be expressed through potential as


"\frac{\partial V}{\partial x}=-F_x,\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space(4)"

In equation (3) either of the pairs of parentheses equal to 0, so from the first parentheses divided by 2 we see it's the total energy equal to zero:


"1/2\cdot m\dot{x}^2+V=0,"

from the second parentheses using equation (4)


"m\ddot{x}=-\frac{\partial V}{\partial x},\\
ma=F_x."

That's Newton's second law.